2. You have a diffraction grating with 2500 lines/cm. You also have a light source that emits light at 2 different wavelengths, 540 nm and 690 nm, at the same time. The screen for your experiment is 1.2 meters from the diffraction grating.

A. What is the line spacing for the grating?

B. What is the difference in the angle of the 2nd bright fringe for each wavelength for this grating?

C. Which wavelength is closer to the center of the diffraction pattern?

D. How would the color separation be different if you used a 1200 lines/cm grating? Explain.

E. What is the width of the central maximum for each wavelength if the same light source is used to illuminate a single slit with a width of 0.045 mm?
Answer #1

(A) The line spacing for the grating is, 2500 lines/cm = 4x104 cm = 4x106m (B) The diffraction angle for 2nd bright fringe foThe difference in the angle of the 2nd bright fringe is, A0=0,- = 20.18° -15.66° = 4.529 (C) From the equation for diffractio(E) The width of central maxima for the wavelength 540 nm is, 22L w= 2(540x109 m)(1.2m) 0.045x10-m = 0.0288m The width of the